# JugglePAC: a Pipelined Accumulation Circuit

### v2

@@ I Introduction
In the realm of modern computing, the ability to efficiently reduce a set of numbers to a single value is fundamental to a wide variety of computational applications. This process, known as reduction, is crucial for tasks ranging from signal processing [1, 2], data compression [3, 4], scientific computing [5, 6], and neural networks [7, 8]. Among the various types of reduction operations, accumulation, which involves summing a dataset to produce a single result, is one of the most essential. As data complexity and scale increase, high-performance accumulation methods that handle large datasets efficiently become more essential.
@@
Input Adder        Input Adder
@@
$a_{0}$ $a_{0}$ 0        $a_{0}$
$a_{1}$ $a_{1}$ 0        $a_{1}$ $a_{0}$ $a_{1}$
$a_{2}$ $a_{2}$ 0        $a_{2}$
$a_{3}$ $a_{0}$ $a_{3}$ $a_{0}$        $a_{3}$ $a_{2}$ $a_{3}$
$a_{4}$ $a_{1}$ $a_{4}$ $a_{1}$        $a_{4}$ $a_{0,1}$
$a_{5}$ $a_{2}$ $a_{5}$ $a_{2}$        $a_{5}$ $a_{4}$ $a_{5}$
Stall $\mathbf{a_{0,3}}$        $b_{0}$ $a_{0,1}$ $a_{2,3}$ $a_{2,3}$
$a_{0,3}$ $a_{1,4}$ $\mathbf{a_{1,4}}$        $b_{1}$ $b_{0}$ $b_{1}$
$\mathbf{a_{2,5}}$        $b_{2}$ $a_{0,1}$ $a_{2,3}$ $a_{4,5}$
       $b_{3}$ $b_{2}$ $b_{3}$ $a_{0:3}$
$a_{0,1,3,4}$ $a_{2,5}$ $a_{0,1,3,4}$        $b_{4}$ $a_{4,5}$ $a_{0:3}$ $b_{0,1}$
$b_{0}$ $b_{0}$ 0        $b_{5}$ $b_{4}$ $b_{5}$
$b_{1}$ $b_{1}$ 0        $b_{6}$ $a_{0:3}$ $a_{4,5}$ $b_{2,3}$
$b_{2}$ $b_{2}$ 0 $\mathbf{a_{0:5}}$        $b_{7}$ $b_{6}$ $b_{7}$ $\mathbf{a_{0:5}}$
Table I depicts an example of a floating-point adder-based accumulator with a pipeline latency of 3 clock cycles, while new input values are fed in every cycle. Using a simple accumulation schedule (SimplePAC), we end up with 3 subsums, 3 cycles after all data inputs are fed to the adder. The subsums are shown as $a_{0,3}$, $a_{1,4}$, and $a_{2,5}$ in bold in Table I. The simplest approach to accumulation would be not to allow a new dataset (called $b_{0:5}$) until the last addition of the current dataset is fed into the adder pipeline (i.e., introduce stalls between consecutive datasets). However, SimplePAC does not accept back-to-back datasets. Using a more intuitive schedule, we can alternate between additions from different sets to maintain a fully pipelined accumulation circuit. Such a schedule is shown in Table I using the JugglePAC schedule, which starts accumulating the data from the consecutive dataset without requiring any stalls.
This work presents JugglePAC, a novel fully pipelined accumulation circuit that optimizes both area and timing using a single floating-point adder. We implement JugglePAC and evaluate its performance across multiple FPGAs, benchmarking it against state-of-the-art solutions. The major contributions of this work are:
@@
    We propose a fully pipelined accumulation circuit, namely JugglePAC, which efficiently handles high-speed, back-to-back variable-length datasets.
@@
    JugglePAC demonstrates improvements over existing solutions by achieving higher frequencies and reducing area complexity.
@@
    Our work introduces a dynamic scheduling mechanism for variable-length datasets, addressing the challenges associated with inefficient control logic in existing designs.
@@
    We implement JugglePAC on two different target FPGAs, specifically the Xilinx XC2VP30 and XC5VLX110T, showing consistent improvements in both area and timing over the state-of-the-art.
@@ II Related Work
Recent work on floating-point accumulation circuits has focused on optimizing area, performance, and complexity. Early designs, such as Luo and Martonosi [9], used carry-save arithmetic and delayed adders for high-performance accumulation but were not fully pipelined, leading to performance bottlenecks due to the required stalls.
Vangal [10] improved on this by introducing a pipelined structure for single-precision floating-point multiply-accumulate operations, yet managing the pipeline complexity remained a challenge. He et al. [11] proposed a group alignment algorithm to improve accuracy but struggled with scalability and efficiency for variable-length datasets.
@@ III JugglePAC
JugglePAC is a novel floating-point accumulation circuit designed to optimize performance and area. This section describes the microarchitecture of JugglePAC and its inter-dataset behavior as well as key challenges such as the ability to mix variable-length datasets.
@@ III-A JugglePAC Microarchitecture
The JugglePAC architecture is built around a floating-point adder, which plays a crucial role in the accumulation process. The fundamental concept of JugglePAC involves scheduling the additions of serially arriving data to achieve efficient accumulation. Given that data typically arrives serially, the adder operates with a throughput of 1, performing additions every 2 cycles. This setup ensures that the adder is utilized 50% of the time, while producing results at least every 2 cycles. Consequently, a single adder is sufficient to maintain pace with the input rate when additions are scheduled effectively. JugglePAC employs a state machine with two distinct states to manage the addition process. In the first state, the inputs are directly added. In the second state, the design handles the addition of any available pair of subsums, as illustrated in Fig. 1. The state machine iterates between these two states every cycle, with an exception for datasets of odd length, where the state machine remains in state 1 for an additional cycle.
To enhance performance for high-throughput applications, JugglePAC is designed to handle back-to-back inputs, thereby avoiding data pile-ups. This approach introduces additional complexity as it necessitates the simultaneous processing of subsums from different datasets. To manage this, each subsum is assigned a unique label, an integer that increments with each new dataset (color-coded with green in Fig. 1). This labeling system is maintained using a shift register with a latency equal to that of the adder as depicted in Fig. 1 and color-coded with purple. The Matching Shift Register block matches subsums from the same dataset and distinguishes between those from different datasets. An additional pipeline stage is added before the inputs of the adder to increase the throughput, which is represented by dashes in the figure.
@@
Scheduling additions of subsums requires efficient control logic. The Pair Identifier (PI) block, shown in Fig. 1 and color-coded with yellow, is responsible for managing this process. The PI receives the results from the adder along with their labels and schedules the additions accordingly. It uses a register for each label, storing incoming subsums and identifying pairs for addition. When a pair is identified, the PI schedules the addition and clears the corresponding register. This control logic allows JugglePAC to juggle between additions from different datasets while minimizing area requirements. The number of registers in the PI depends on the label size, with L representing the maximum number of labels. To handle situations where the adder may not always be available for scheduled additions, JugglePAC employs a FIFO buffer to temporarily store ready-to-add subsum pairs along with their labels. The FIFO is read every 2 cycles, with a maximum depth of $\lceil\log(p)\rceil$, where $p$ is the adder’s latency. The FIFO’s design, which utilizes registers, ensures efficient data management.
@@
The state machine within JugglePAC manages the accumulation process by maintaining a low area and timing complexity. It alternates between adding serial input data and processing data from the FIFO. A sample scheduling scenario using the JugglePAC approach is illustrated in Table I, demonstrating how the accumulation of dataset $b_{0:N}$ begins before the results of dataset $a_{0:5}$ are completed. The simple control logic for the state machine and pair identification allows JugglePAC to maintain a low area complexity and critical path, outperforming the state-of-the-art.
For output identification, JugglePAC uses a counter to track the number of additions performed. The counter increments with each addition from state 1 and decrements with each addition from state 0. The system skips incrementing on the first addition of state 1, ensuring the counter returns to zero after each operation. The counter reaches a maximum value of $\lceil\frac{p+2}{2}\rceil$, where $p$ is the adder’s latency, and returns to zero upon completing the final addition. Separate counters are used for each simultaneous accumulation, with $L$ representing the maximum number of counters. The output identifier is represented in Fig. 1 and color-coded with red.
@@ III-B Inter-Dataset Behavior
The JugglePAC architecture allows for the parallel processing of datasets, with the label size determining the maximum number of datasets that can be handled simultaneously. A smaller label size reduces area complexity but introduces a minimum dataset length requirement. When the number of datasets exceeds $L$, the circuit may mix accumulations from different datasets, leading to incorrect results. Therefore, the minimum dataset length is identified through comprehensive testing with variable-length datasets.
Increasing the label size enhances the circuit’s capability to handle more datasets but also increases the logical resources required for the PI and output identification modules. This tradeoff between area complexity and minimum dataset length is summarized in Section IV, which shows the area complexity increasing with larger label sizes.
JugglePAC’s dynamic accumulation approach results in latency that depends on the current and previous dataset lengths. For label sizes less than 3, the circuit maintains consistent latency, ensuring that the results are produced in input order. However, with larger label sizes, the minimum dataset length can lead to results being output in an order that deviates from the input sequence. Additional control logic can be used to reorder results based on the label or to output the label itself, allowing the system to identify the dataset to which the result belongs. Setting the minimum dataset length to 19 ensures that JugglePAC consistently produces ordered results. This design achieves a balance between high frequency and low area complexity, surpassing existing state-of-the-art solutions.
Figure 1: JugglePAC microarchitecture features a Floating-point Adder and key components including a state machine for managing additions, a Matching Shift Register for labeling subsums, and a Pair Identifier block for scheduling. A FIFO buffer ensures efficient data handling and prevents pile-ups, while an Output Identifier tracks the number of additions performed. This design optimizes throughput and minimizes area complexity.
@@ IV Implementation Results
We evaluated the JugglePAC floating-point accumulation circuit against existing designs, focusing on area complexity, latency, throughput, and frequency. Table II summarizes these metrics for JugglePAC and other designs.
@@
In terms of latency, JugglePAC performs competitively. For a label size of 2 and minimum dataset length of 22, JugglePAC achieves a latency of approximately 1.077 $\mu$s, comparable to or better than most previous designs such as DSA and SSA [13]. Its throughput is also high, outperforming designs like DB [19] and BTTP [20], especially with larger datasets.
JugglePAC operates at a frequency of 208 MHz, exceeding many previous designs like FPACC [16] and FCBT [13]. It achieves the lowest ”Slices ×$\times\mu$s” score, reflecting superior efficiency in balancing area and performance.
The full pipelining of JugglePAC enhances its performance and resource utilization, making it effective in high-speed operations. However, using a single floating-point adder might limit performance in cases requiring multiple adders. JugglePAC’s minimum dataset length varies with label size, ensuring accurate accumulation but potentially limiting flexibility. Future work could explore the incorporation of multiple adders to address these limitations and enhance performance.
Overall, JugglePAC represents a significant advancement in floating-point accumulation circuits, with strong performance across key metrics. The results highlight its potential for various applications and provide a basis for future research and optimization.
Design Label Min. Dataset Adders Slices BRAMs Frequency Total Latency Slices×$\times\mu$s FPGA
Size Length (MHz) clock cycles $\mu$s
MFPA [14] - - 4 4,991 2 207 198 0.957 4,776 XC2VP30
AeMFPA [14] - - 2 3,130 14 204 198 0.970 3,036
$Ae^{2}$ [14] - - 2 3,737 2 144 198 1.370 5,120
FAAC [15] - - 3 6,252 0 199 176 1.086 6,790
FCBT [13] - - 2 2,859 10 170 $\leq$ 475 $\leq$ 2.794 7,988
DSA [13] - - 2 2,215 3 142 232 1.634 3,619
SSA [13] - - 1 1,804 6 165 $\leq$ 520 $\leq$ 3.152 5,686
DB [19] - - 1 1,749 6 188 $\leq$ 199 $\leq$ 1.058 1850
JugglePAC 1 74 1 1,439 0 208 $\leq$ 220 $\leq$ 1.058 1,522
JugglePAC 2 22 1 1,796 0 208 $\leq$ 224 $\leq$ 1.077 1,934
JugglePAC 3 10 1 2,343 0 208 $\leq$ 224 $\leq$ 1.077 2,523
FPACC [16] - - - 683 - 247 - - - VC5VSX50T
BTTP [20] - - 1 648 9.5 305 - - - XC5VLX110T
JugglePAC 2 22 1 578 0 334 $\leq$224 $\leq$ 0.671 388

### v3

@@ I Introduction
In the realm of modern computing, the ability to efficiently reduce a set of numbers to a single value is fundamental to a wide variety of computational applications. This process, known as reduction, is crucial for tasks ranging from signal processing [1, 2], data compression [3, 4], scientific computing [5, 6], and neural networks [7, 8]. Among reduction operations, accumulation, which sums a dataset into a single result, is one of the most fundamental. As data complexity and scale increase, high-performance accumulation methods that handle large datasets efficiently become essential.
@@
Cycle Input Adder     Input Adder
@@
0 $a_{0}$ $a_{0}$ 0    $a_{0}$
1 $a_{1}$ $a_{1}$ 0    $a_{1}$ $a_{0}$ $a_{1}$
2 $a_{2}$ $a_{2}$ 0    $a_{2}$
3 $a_{3}$ $a_{3}$ $a_{0}$ $a_{0}$   $a_{3}$ $a_{2}$ $a_{3}$
4 $a_{4}$ $a_{4}$ $a_{1}$ $a_{1}$   $a_{4}$ $a_{0,1}$
5 $a_{5}$ $a_{5}$ $a_{2}$ $a_{2}$   $a_{5}$ $a_{4}$ $a_{5}$
6 Stall $a_{0,3}$   $b_{0}$ $a_{0,1}$ $a_{2,3}$ $a_{2,3}$
7 $a_{0,3}$ $a_{1,4}$ $a_{1,4}$   $b_{1}$ $b_{0}$ $b_{1}$
8 $a_{2,5}$   $b_{2}$ $a_{4,5}$
9    $b_{3}$ $b_{2}$ $b_{3}$ $a_{0:3}$
10 $a_{2,5}$ $a_{0,1,3,4}$ $a_{0,1,3,4}$   $b_{4}$ $a_{4,5}$ $a_{0:3}$ $b_{0,1}$
11 $b_{0}$ $b_{0}$ 0    $b_{5}$ $b_{4}$ $b_{5}$
12 $b_{1}$ $b_{1}$ 0    $b_{6}$ $b_{0,1}$ $b_{2,3}$ $b_{2,3}$
13 $b_{2}$ $b_{2}$ 0 $\mathbf{a_{0:5}}$   $b_{7}$ $b_{6}$ $b_{7}$ $\mathbf{a_{0:5}}$
Table I depicts an example of a floating-point accumulator using a single adder with a pipeline latency of 3 clock cycles, where new input values are fed in each cycle. Different datasets are represented by distinct letters, such as $a$ and $b$, with subscripts indicating the input indexes. For example, $a_{0}$ refers to the first element in dataset $a$, and $a_{0,1}$ denotes the sum $a_{0}+a_{1}$.
In an accumulation circuit, the adder’s output feeds back to the corresponding new input. For SimplePAC, due to the adder’s pipeline latency of 3 clock cycles and the unavailability of immediate outputs, the first 3 inputs are added with zeros to initialize the accumulation. Once the outputs (subsums) are ready, they are summed with the next input, either from the fed input or previous subsums. In contrast, JugglePAC eliminates the need for zero-padding during the initial stage. Instead, when the first and second inputs are ready, they are immediately summed together. For example, in cycle 1, inputs $a_{0}$ and $a_{1}$ are summed without any zero-padding.
In SimplePAC, subsums are accumulated using a fixed reduction tree with P variables, where P is the adder’s pipeline depth. However, this approach does not handle consecutive datasets properly, as results from datasets $a$ and $b$ may mix, leading to incorrect outcomes. While introducing stalls between datasets prevents this, JugglePAC improves upon this by alternating between datasets, ensuring the accumulation circuit stays fully pipelined without any stalls, as shown in Table I.
This work presents JugglePAC, a novel fully pipelined accumulation circuit that optimizes both performance and area-timing efficiency using a single floating-point adder. The major contributions of this work are:
@@ 13 $b_{2}$ $b_{2}$ 0 $\mathbf{a_{0:5}}$   $b_{7}$ $b_{6}$ $b_{7}$ $\mathbf{a_{0:5}}$
    We propose JugglePAC, a fully pipelined accumulation circuit with a novel dynamic scheduling mechanism that efficiently handles high-speed, back-to-back variable-length datasets, overcoming control logic challenges and improving both performance and area efficiency.
@@
    Our design introduces innovative operation scheduling and dataset identification using labels, simplifying control logic, reducing area, and improving critical path performance in supported applications.
@@
    We implement JugglePAC across multiple target FPGAs, specifically the Xilinx XC2VP30 and XC5VLX50T. JugglePAC achieves the highest clock frequency for label widths 1, 2, and 3, reaching 208 MHz for the XC2VP30 and 334 MHz for the XC5VLX50T, outperforming existing solutions.
@@
[…]
@@ II Related Work
Recent work on floating-point accumulation circuits has aimed to optimize area, performance, and complexity. Early designs, such as Luo and Martonosi [9], used carry-save arithmetic and delayed adders but lacked full pipelining, causing performance stalls. Vangal [10] introduced pipelined multiply-accumulate units, though managing pipeline complexity remained challenging. He et al. [11] improved accuracy using a group alignment algorithm but struggled with scalability for variable-length datasets.
@@ III JugglePAC
JugglePAC is a novel floating-point accumulator optimized for performance and area. This section details its microarchitecture, inter-dataset behavior, and challenges like mixing variable-length datasets.
@@ III-A JugglePAC Microarchitecture
Figure 1: JugglePAC’s microarchitecture includes a Floating-Point Adder and key components such as a state machine for addition management, a Matching Shift Register for subsum labeling, and a Pair Identifier for scheduling. A FIFO buffer ensures smooth data flow, preventing congestion, while an Output Identifier tracks addition counts. This design enhances throughput while minimizing area complexity.
@@
The JugglePAC architecture is centered around a floating-point adder, which is critical for the accumulation process. It schedules serially arriving data to ensure efficient addition, operating at a throughput of 1 and performing additions every 2 cycles, resulting in 50% utilization of the adder. This setup allows a single adder to keep up with the input rate when scheduling is managed effectively. JugglePAC employs a state machine with two states: in the state 1, inputs are directly added, while in state 0, available subsums are accumulated, as illustrated in Fig. 1. The state machine alternates between these states each cycle, except for datasets with an odd length, where it remains in state 1 for an additional cycle.
To enhance performance in high-throughput applications, JugglePAC processes back-to-back inputs, preventing data pile-ups. This requires handling subsums from multiple datasets simultaneously. Each subsum is assigned a unique incrementing label (green in Fig. 1) to differentiate datasets. A shift register, with latency matching the adder (purple in Fig. 1), maintains this labeling. The Matching Shift Register ensures subsums are correctly paired within the same dataset while distinguishing those from different datasets. To enhance throughput, additional pipeline stages are introduced before the adder inputs. The registers corresponding to these pipeline stages are depicted by short thick lines in Fig. 1, placed along the connections between different blocks. Depending on the desired application, intentional gaps or delays may occur within or between datasets, during which the valid input is set to 0. In these cases, the green multiplexer in Fig. 1 forces the input to 0 during accumulation, ensuring the output remains unaffected.
@@
Scheduling the additions of subsums requires efficient control logic. The Pair Identifier (PI) block, shown in Fig. 1 and highlighted in yellow, manages this process. It stores incoming subsums in registers, identifies addition pairs, and schedules the operations accordingly. The number of registers depends on the label size (L). To handle periods when the adder is unavailable, JugglePAC uses a FIFO buffer to store ready-to-add subsum pairs along with their labels. This approach allows efficient juggling between datasets while minimizing area requirements. After the last input is received, the circuit holds at most $p$ subsums in the pipeline, which require $\log(p)$ cycles to reduce. Thus, the next accumulation needs to buffer at most log(p) values. Consequently, the FIFO has a depth of $\lceil\log(p)\rceil$, where $p$ is the adder’s latency. The FIFO’s design, which utilizes registers, ensures efficient data management.
@@
The state machine in JugglePAC controls the accumulation process with low area and timing complexity. It alternates between adding serial input data and processing FIFO data, managed by the purple-colored multiplexers. Table I shows a sample schedule, where accumulation of dataset $b_{0:N}$ begins before the completion of $a_{0:5}$. The simple state machine and pair identification logic help JugglePAC achieve low area and critical path delays, outperforming state-of-the-art designs
The simplest way to identify when a dataset’s final output is obtained is by counting inputs and additions. For $N$ values, N1N-1 additions are required; however, $N$ is not fixed. An up-down counter handles this by incrementing for each input and decrementing for each addition. JugglePAC optimizes this by tracking the number of additions performed from each state. The output identifier is shown in Fig. 1, color-coded in red. The counter increments for each addition from state 1 and decrements for each addition from state 0, skipping the first addition of state 1 to reset after each operation.
For the first $p+3+(1-(p\bmod 2))$ cycles (where $p$ is the adder’s latency), only state 1 performs additions. The term $p+3$ accounts for the cycles required to generate two subsums, while the additional $1-(p\bmod 2)$ reflects the scheduling delay before state 0 can perform its first addition. After this period, state 1 continues to perform additions more frequently than state 0 until a steady state is reached (when both states perform additions at the same rate). The counter reaches a maximum of $p+2$, representing the buffer size before state 0 schedules its next addition (i.e., before the FIFO). The additional 2 cycles in this value account for the buffers in the Pair Identifier block and a pipeline register between the FIFO and the adder.
@@ III-B Inter-Dataset Behavior
The JugglePAC architecture allows for parallel processing of datasets, with the label size determining the maximum number of datasets that can be handled simultaneously. When the number of datasets exceeds $2^{L}$, different datasets may be assigned the same label, causing mixing and incorrect results. However, if the dataset lengths are long enough, this issue is avoided. To ensure this, JugglePAC introduces a minimum dataset length constraint, which depends on the label size.
To calculate the minimum dataset length, we consider the maximum overlap between consecutive datasets, which defines this lower bound. Starting from the simple case where the label size is 1, only two datasets can overlap. Each dataset can have at most $p$ subsums in the system after its last input is received, where $p$ is the adder’s latency, resulting in $\lceil\log p\rceil$ levels in the binary reduction tree. The output identification logic excludes the initial addition from state 1, which effectively increases the scheduling latency by an additional $p$ cycles. Additionally, JugglePAC requires datasets to be larger than 4 to be recognized due to the output identification logic, which is accounted for as a constant offset and a minimum. As the label size increases, more overlap is possible. Using the base case, the minimum dataset length is calculated using (1).
$\max(\left\lceil\frac{(1+\lceil\log p\rceil)p+4}{2^{L}-1}\right\rceil,4)$ (1)
JugglePAC’s dynamic accumulation approach leads to latency variations based on the current and previous dataset lengths. For label widths less than 3, latency stays consistent, with variations under 5 cycles, maintaining input order. For larger widths, short datasets can cause results to appear out of order. This can be corrected using control logic to reorder outputs based on the label or by outputting the label itself. Our analysis shows that setting the minimum dataset length to 19 consistently preserves input order across all test cases. This design effectively balances high frequency and low area complexity, outperforming current state-of-the-art solutions.
@@ IV Implementation Results
We evaluated the JugglePAC floating-point accumulation circuit against prior designs, focusing on area, latency, throughput, and frequency. Table II summarizes these metrics, with dashes (–) indicating unavailable or inapplicable values. Notably, the label width parameter is unique to JugglePAC; since the minimum dataset length depends on it, this metric is not relevant to previous works.
[…]
MFPA [14] - - 4 4,991 - - 2 207 198 0.957 4,776 XC2VP30
AeMFPA [14] - - 2 3,130 - - 14 204 198 0.970 3,036
$Ae^{2}$MFPA [14] - - 2 3,737 - - 2 144 198 1.370 5,120
FAAC [15] - - 3 6,252 - - 0 199 176 1.086 6,790
FCBT [13] - - 2 2,859 - - 10 170 $\leq$ 475 $\leq$ 2.794 7,988
DSA [13] - - 2 2,215 - - 3 142 232 1.634 3,619
SSA [13] - - 1 1,804 - - 6 165 $\leq$ 520 $\leq$ 3.152 5,686
DB [19] - - 1 1,749 - - 6 188 $\leq$ 199 $\leq$ 1.058 1850
JugglePAC 1 74 1 1,439 1763 2010 0 208 $\leq$ 220 $\leq$ 1.058 1,522
JugglePAC 2 25 1 1,796 1910 2592 0 208 $\leq$ 224 $\leq$ 1.077 1,934
JugglePAC 3 11 1 2,343 2199 3418 0 208 $\leq$ 224 $\leq$ 1.077 2,523
JugglePAC 4 5 1 3,339 2800 5086 0 159 $\leq$ 224 $\leq$ 1.409 4,704
FPACC [16] - - - 683 1744 2234 - 247 - - - VC5VSX50T
JugglePAC 2 25 1 625 1702 1607 0 334 $\leq$224 $\leq$ 0.671 419 XC5VLX50T
BTTP [20] - - 1 648 1554 1495 9.5 305 - - - XC5VLX110T
@@
[…]
JugglePAC’s full pipelining improves performance and resource utilization, excelling in high-speed operations. However, using a single floating-point adder may limit performance in multi-adder scenarios. Its minimum dataset length varies with label width, ensuring accurate accumulation but potentially limiting flexibility. Future work could explore using multiple adders to overcome these limitations and boost performance.